# Multi-marginal temporal Schrödinger Bridge Matching for video generation from unpaired data

### v1

@@ Abstract
Many natural dynamic processes -such as in vivo cellular differentiation or disease progression- can only be observed through the lens of static sample snapshots. While challenging, reconstructing their temporal evolution to decipher underlying dynamic properties is of major interest to scientific research. Existing approaches enable data transport along a temporal axis but are poorly scalable in high dimension and require restrictive assumptions to be met. To address these issues, we propose Multi-Marginal temporal Schrödinger Bridge Matching (MMtSBM) for video generation from unpaired data, extending the theoretical guarantees and empirical efficiency of Diffusion Schrödinger Bridge Matching (Shi et al., 2023) by deriving the Iterative Markovian Fitting algorithm to multiple marginals in a novel factorized fashion. Experiments show that MMtSBM retains theoretical properties on toy examples, achieves state-of-the-art performance on real world datasets such as transcriptomic trajectory inference in 100 dimensions, and for the first time recovers couplings and dynamics in very high dimensional image settings. Our work establishes multi-marginal Schrödinger bridges as a practical and principled approach for recovering hidden dynamics from static data.111code: github.com/ICLRMMtDSBM/MMDSBM_ILCR $|$ website: mmdsbm.notion.site
@@ 1 Introduction
The modelization of this problem is inherently probabilistic, given both the variability occurring in complex natural processes and the uncertainty of the observation. We thus ask the question: ”What is the most probable evolution of an existing data point, given uncoupled samples of the same process acquired across different times?”. This point of view has notably been developed in the Schrödinger Bridge (SB) theory (Schrödinger, 1931). The SB is the unique stochastic process whose marginals at start and end times match given probability distributions while minimizing the Kullback–Leibler (KL) divergence w.r.t. a given reference process. The SB also happens to solve a regularized Optimal Transport (OT) problem (Léonard, 2014). Recent major advances in statistical learning of SBs have allowed using this framework between complex empirical distributions (De Bortoli et al., 2021; Wang et al., 2021), achieved important improvements in their efficiency (Shi et al., 2023; Bortoli et al., 2024), extended it to the multi-marginal setting and explored various additional constraints such as smooth trajectories (Chen et al., 2023a; Hong et al., 2025), and spline-valued trajectories (Theodoropoulos et al., 2025). A few methods have been proposed to solve the SB problem in an applied machine learning setting. (De Bortoli et al., 2021) uses iterative proportional fitting (IPF) (Kullback, 1968), the general continuous analogue of the well known Sinkhorn algorithm (Cuturi, 2013). Subsequent works have explored alternative training schemes based on likelihood bounds (Chen et al., 2023b) or on the dual algorithm of IPF: Iterative Markovian Fitting (IMF) (Shi et al., 2023).
@@
Contributions This paper makes the following contributions: 1) We define the multi-marginal temporal Schrödinger Bridge problem and demonstrate its fundamental properties 2) We introduce a novel factorized extension of the IMF algorithm presented in Shi et al. (2023) to multiple iterative marginals in an efficient and principled way. 3) We produce a convergence analysis of the algorithm under asymptotic hypotheses. 4) We demonstrate the soundness of the method on low-to-medium-dimensional examples, achieving state-of-the-art results on the widely reported Embryoid Body (EB) dataset (Moon et al., 2019). 5) We scale up to 7 iterative marginals in a very high-dimensional image setting, presenting for the first time a coherent video generation algorithm from purely time-unpaired data samples.
@@
We adopt the notations from Shi et al. (2023). We denote by $\mathcal{P}(C)$ the space of path measures, with $\mathcal{P}(C)=\mathcal{P}(C([0,T],\mathbb{R}^{d}))$, where $C([0,T],\mathbb{R}^{d})$ is the space of continuous functions from $[0,T]$ to $\mathbb{R}^{d}$. The subset of Markov path measures associated with the diffusion $dX_{t}=v_{t}(X_{t})dt+\sigma_{t}dB_{t}$, with $\sigma,v$ locally Lipschitz, is denoted $\mathcal{M}$. For $\mathbb{Q}$ induced by $(\sqrt{\varepsilon}B_{t})_{t\in[0,1]}$, $\varepsilon>0$, where $(B_{t})_{t\geq 0}$ is a $d$-dimensional Brownian motion, the reciprocal class of $\mathbb{Q}$ is $\mathcal{R}(\mathbb{Q})$ (see Def 3.2) For $\mathbb{P}\in\mathcal{P}(C)$, we denote by $\mathbb{P}_{t}$ its marginal at time $t$, by $\mathbb{P}_{s,t}$ the joint law at times $s,t$, and by $\mathbb{P}_{s|t}$ the conditional law at $s$ given $t$. We write $\mathbb{P}_{|t_{i},t_{j}}\in\mathcal{P}(C)$ for the path distribution on $(t_{i},t_{j})$ given the endpoints $t_{i}$ and $t_{j}$; e.g., $\mathbb{Q}_{|t_{i},t_{j}}$ is a scaled Brownian bridge. Unless otherwise specified, $\nabla$ refers to gradients w.r.t. $x_{t}$ at time $t$. For a joint law $\Pi_{0,T}$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, the mixture of bridges measure is $\Pi=\Pi_{0,T}\mathbb{P}_{|0,T}\in\mathcal{P}(C)$ with $\Pi(\cdot)=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\mathbb{P}_{|0,T}(\cdot|x_{0},x_{T})d\Pi_{0,T}(x_{0},x_{T})$. The entropy of a process w.r.t. the Brownian motion is denoted $\mathcal{H}$. Finally, for $\pi_{0},\pi_{T}\in\mathcal{P}(X)$, the Kullback–Leibler divergence is $\mathrm{KL}(\pi_{0}\|\pi_{T})=\int_{X}\log\!\left(\tfrac{d\pi_{0}}{d\pi_{T}}(x)\right)\,d\pi_{0}(x)$.
@@ 2.1 The Schrödinger Bridge problem
The Schrödinger Bridge problem (Schrödinger, 1931) seeks the most likely stochastic evolution between marginals $\mu_{0},\mu_{T}$ under a reference law $\mathbb{Q}$. It admits both a dynamic formulation,
@@ 2.2 Iterative Markovian Fitting (IMF)
The SB solution is the unique path measure that is both Markovian and belongs to the reciprocal class of $\mathbb{Q}$ while matching marginals (Léonard, 2014). This motivates the Iterative Markovian Fitting (IMF) algorithm (Shi et al., 2023; Peluchetti, 2023), which alternates between reciprocal and Markov projections:
$\mathbb{P}^{2n+1}=\operatorname{proj}_{\mathcal{M}}(\mathbb{P}^{2n}),\qquad\mathbb{P}^{2n+2}=\operatorname{proj}_{\mathcal{R}(\mathbb{Q})}(\mathbb{P}^{2n+1}).$ (3)
These projections admit KL variational characterisations (App. B), and the iteration converges to $\mathbb{P}^{\star}$.
In practice, IMF is implemented by learning the drift of the Markovian projection via a bridge-matching loss (see App. B). Compared to Iterative Proportional Fitting (IPF), IMF preserves both marginals simultaneously and is more efficient (details in App. B).
@@ 3 Multi-marginal temporal Schrödinger Bridge Matching
All proofs can be found in the Appendix B.5.
@@ 3.1 Multi-Marginal temporal Schrödinger Bridge Problem
Let $0=t_{0}<t_{1}<\cdots<t_{N}=T$ be a fixed time grid, and let $\mu_{0},...,\mu_{k},...,\mu_{T}\in\mathcal{P}(\mathbb{R}^{d})$ denote prescribed marginals at times $(t_{k})_{k=0,\dots,N}$. Given a reference process $\mathbb{Q}$ on $C([0,T],\mathbb{R}^{d})$, the multi-marginal Schrödinger Bridge problem (MMSB) is defined as
$\mathbb{P}^{\star}\;=\;\underset{\mathbb{P}\in\mathcal{P}(C)}{\operatorname{argmin}}\,\mathrm{KL}\!\left(\mathbb{P}\,\|\,\mathbb{Q}\right)\quad\text{subject to}\quad X_{t_{k}}\sim\mu_{k},\;\;k=0,\dots,N$ (4)
Connection to Multi-marginal Optimal Transport
If $\mathbb{Q}$ is associated with a Brownian motion, the induced reference coupling $\mathbb{Q}_{t_{0},\dots,t_{N}}$ is characterized by independent Gaussian increments $X_{t_{i+1}}-X_{t_{i}}\sim\mathcal{N}(0,\sigma^{2}(t_{i+1}-t_{i}))$. By evaluating the KL term, 4 can therefore be rewritten as
$\Pi^{\star}=\arg\min_{\Pi\in\mathcal{P}((\mathbb{R}^{d})^{N})}\Bigg\{\mathbb{E}_{X\sim\Pi}\Bigg[\sum_{i=0}^{N-1}\frac{1}{t_{i+1}-t_{i}}\,\|X_{t_{i+1}}-X_{t_{i}}\|^{2}\Bigg]-2\sigma^{2}T\mathcal{H}(\Pi)\;\;:\;\;\Pi_{i}=\mu_{t_{i}},\;\forall i\Bigg\}$
This is precisely an entropy-regularised multi-marginal OT problem with a time-structured quadratic cost $c(x_{0},\dots,x_{N})=\sum_{i=0}^{N-1}\frac{1}{t_{i+1}-t_{i}}\|x_{i+1}-x_{i}\|^{2}$ and entropy-regularisation parameter $\varepsilon=2\sigma^{2}$. [^2] This formulation is particularly interesting in practice: in many applications we only observe static snapshots $\mu_{t_{i}}$ without any coupling between them, and the multi-marginal Schrödinger Bridge not only provides a consistent path modeling but also induces the most natural coupling between the observed distributions.
@@
These properties ensure that we can restrict our search to Markovian (and therefore reciprocal B.1) measures, which will be central to the projection algorithms introduced later. Finally, the explicit form of the solution further clarifies its structure.
@@
Under mild assumptions,
@@ 3.2.1 Multi-marginal Markov and Reciprocal projections
By the Markovian projection theorem of Gyöngy (1986), and as further developed in Peluchetti (2023); De Bortoli et al. (2021), the process $M^{\star}$ is Markov and matches the one-dimensional marginals of the original factorized law $\Pi$.
@@
Assume that $\sigma_{t}>0$. Let $M^{\star}=\operatorname{proj}_{\mathcal{M}}(\Pi)$ be the Markovian projection of $\Pi$ as in Definition 3.3. Then we work under the assumption that the variational characterization still holds in the factorized setting ((Shi et al., 2023; Gyöngy, 1986; Peluchetti, 2023)), i.e.
@@
$KL(\Pi\,\|\,M^{\star})=\frac{1}{2}\int_{0}^{T}\mathbb{E}_{\Pi_{t_{i},t}}\!\left[\frac{1}{\sigma_{t}^{2}}\Big\|\sigma_{t}^{2}\,\mathbb{E}_{\Pi_{t_{i+1}\mid t}}\!\big[\nabla\log\mathbb{Q}^{\,|t_{i},t_{i+1}}_{t}(X_{t_{i+1}}\mid X_{t})\,\big|\,X_{t}\big]-v_{t}^{\star}(X_{t})\Big\|^{2}\right]dt$
@@
Conjecture 3.1 (Analogue of Léonard (2014) Theorem 2.12).
@@ 3.2.2 Iterative Markovian Factorized Fitting
$KL(P^{n+1}\,\|\,P^{\star})\;\leq\;KL(P^{n}\,\|\,P^{\star})<\infty,\qquad\lim_{n\to\infty}KL(P^{n}\,\|\,P^{\star})=0$
@@
Under mild assumptions, the IMFF sequence $(\mathbb{P}^{n})_{n\in\mathbb{N}}$ admits at least one fixed point $\mathbb{P}^{\star}$, and we have
@@
Moreover, denoting by $\mathbb{P}^{\mathrm{MMSB}}$ the true multi-marginal Schrödinger bridge and by $\mathbb{P}^{\mathrm{pair}}$ the classical Schrödinger bridge constructed pairwise, the limiting solution satisfies the inequality
$KL(\mathbb{P}^{\mathrm{MMSB}}\,\|\,\mathbb{Q})\;\;\leq\;\;KL(\mathbb{P}^{\star}\,\|\,\mathbb{Q})\;\;\leq\;\;KL(\mathbb{P}^{\mathrm{pair}}\,\|\,\mathbb{Q})$
where $\mathbb{Q}$ is the chosen reference process. Thus, $\mathbb{P}^{\star}$ can be interpreted as an approximate multi-marginal Schrödinger Bridge, lying between the optimal multi-marginal solution and the pairwise construction.

### v2

# Multi-marginal temporal Schrödinger Bridge Matching from unpaired data

@@ Abstract
Many natural dynamic processes -such as in vivo cellular differentiation or disease progression- can only be observed through the lens of static sample snapshots. While challenging, reconstructing their temporal evolution to decipher underlying dynamic properties is of major interest to scientific research. Existing approaches enable data transport along a temporal axis but are poorly scalable in high dimension and require restrictive assumptions to be met. To address these issues, we propose Multi-Marginal temporal Schrödinger Bridge Matching (MMtSBM) for video generation from unpaired data, extending the theoretical guarantees and empirical efficiency of Diffusion Schrödinger Bridge Matching (shi2023diffusionschrodingerbridgematching) by deriving the Iterative Markovian Fitting algorithm to multiple marginals in a novel factorized fashion. Experiments show that MMtSBM retains theoretical properties on toy examples, achieves state-of-the-art performance on real world datasets such as transcriptomic trajectory inference in 100 dimensions, and for the first time recovers couplings and dynamics in very high dimensional image settings, effectively generating temporally coherent videos from purely unpaired data. Our work establishes multi-marginal Schrödinger bridges as a practical and theoretically principled approach for recovering hidden dynamics from static data.
code: github.com/ICLRMMtDSBM/MMDSBM_ILCR $|$ website: mmdsbm.notion.site
@@ 1 Introduction
The modelization of this problem is inherently probabilistic, given both the variability occurring in complex natural processes and the uncertainty of the observation. We thus ask the question: ”What is the most probable evolution of an existing data point, given uncoupled samples of the same process acquired across different times?”. This point of view has notably been developed in the Schrödinger Bridge (SB) theory (schrodinger1931). The SB is the unique stochastic process whose marginals at start and end times match given probability distributions while minimizing the Kullback–Leibler (KL) divergence w.r.t. a given reference process. The SB also happens to solve a regularized Optimal Transport (OT) problem (leonard2014survey). Some recent works such as mmsb; lavenant have explored the theoretical setting of multiple marginals. Recent major advances in statistical learning of SBs have allowed using this framework between complex empirical distributions (debortoli2021diffusionschrodingerbridgeapplications; wang2021deepgenerativelearningschrodinger), achieved important improvements in their efficiency (shi2023diffusionschrodingerbridgematching; debortoli2024schrodingerbridgeflowunpaired), extended it to the multi-marginal setting and explored various additional constraints such as smooth trajectories (chen2023deepmomentummultimarginalschrodinger; hong2025trajectoryinferencesmoothschrodinger), and spline-valued trajectories (theodoropoulos2025momentummultimarginalschrodingerbridge). A few methods have been proposed to solve the SB problem in an applied machine learning setting. debortoli2021diffusionschrodingerbridgeapplications use iterative proportional fitting (IPF) (kullback), the general continuous analogue of the renown Sinkhorn algorithm (cuturi2013sinkhorn). Subsequent works have explored alternative training schemes based on likelihood bounds (chen2023likelihoodtrainingschrodingerbridge) or on the dual algorithm of IPF: Iterative Markovian Fitting (IMF) (shi2023diffusionschrodingerbridgematching).
A closely related line of work is flow matching (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning; albergo2023buildingnormalizingflowsstochastic). These methods have explored OT variants since their inception and have been extended to the multi-marginal setting as well as connected to the Schrödinger Bridge theory (tong2024improvinggeneralizingflowbasedgenerative; tong2024simulationfreeschrodingerbridgesscore; kapuśniak2024metricflowmatchingsmooth).
Concurrent to our work is park2025multimarginalschrodingerbridgematching; we note that they do not scale to video experiments.
@@
Contributions
This paper makes the following contributions:
  1. 1.
    We define the multi-marginal temporal Schrödinger Bridge problem and demonstrate its fundamental properties (existence and uniqueness of the solution).
  2. 2.
    We introduce a novel factorized extension of the IMF algorithm presented in shi2023diffusionschrodingerbridgematching to multiple iterative marginals in a way that is efficient –because parallelized along times, and principled –because mathematical sound and with a concrete algorithm closely following theory.
  3. 3.
    We produce a convergence analysis of the algorithm under asymptotic hypotheses.
  4. 4.
    We demonstrate the soundness of the method on low-to-medium-dimensional examples, and achieve state-of-the-art results against comparable methods on $2$ widely reported single-cell transcriptomic benchmarks (Moon2019; open-problems-multimodal).
  5. 5.
    We scale up to 7 iterative marginals in a very high-dimensional image setting, presenting for the first time a coherent video generation algorithm from purely time-unpaired data samples.
@@
We adopt the notations from shi2023diffusionschrodingerbridgematching. We denote by $\mathcal{P}(C)$ the space of path measures, with $\mathcal{P}(C)=\mathcal{P}(C([0,T],\mathbb{R}^{d}))$, where $C([0,T],\mathbb{R}^{d})$ is the space of continuous functions from $[0,T]$ to $\mathbb{R}^{d}$. The subset of Markov path measures associated with the diffusion $dX_{t}=v_{t}(X_{t})dt+\sigma_{t}dB_{t}$, with $\sigma,v$ locally Lipschitz, is denoted $\mathcal{M}$. We denote $(B_{t})_{t\geq 0}$ the $d$-dimensional Brownian motion. For a process $\mathbb{Q}$, the reciprocal class of $\mathbb{Q}$ is $\mathcal{R}(\mathbb{Q})$. For $\mathbb{P}\in\mathcal{P}(C)$, we denote by $\mathbb{P}_{t}$ its marginal at time $t$, by $\mathbb{P}_{s,t}$ the joint law at times $s,t$, and by $\mathbb{P}_{s|t}$ the conditional law at $s$ given $t$. We write $\mathbb{P}_{|t_{i},t_{j}}\in\mathcal{P}(C)$ for the path distribution on $(t_{i},t_{j})$ given the endpoints $t_{i}$ and $t_{j}$; e.g., $\mathbb{Q}_{|t_{i},t_{j}}$ is a scaled Brownian bridge. Unless otherwise specified, $\nabla$ refers to gradients w.r.t. $x_{t}$ at time $t$. For a joint law $\Pi_{0,T}$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, the mixture of bridges measure is $\Pi=\Pi_{0,T}\mathbb{P}_{|0,T}\in\mathcal{P}(C)$ with $\Pi(\cdot)=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\mathbb{P}_{|0,T}(\cdot|x_{0},x_{T})d\Pi_{0,T}(x_{0},x_{T})$. The entropy of a process w.r.t. the Brownian motion is denoted $\mathcal{H}$. Finally, for $\pi_{0},\pi_{T}\in\mathcal{P}(X)$, the Kullback–Leibler divergence is $\mathrm{KL}(\pi_{0}\|\pi_{T})=\int_{X}\log\!\left(\tfrac{d\pi_{0}}{d\pi_{T}}(x)\right)\,d\pi_{0}(x)$.
@@ 2.1 The Schrödinger Bridge problem
The Schrödinger Bridge problem (schrodinger1931) seeks the most likely stochastic evolution between marginals $\mu_{0},\mu_{T}$ under a reference law $\mathbb{Q}$. It admits both a dynamic formulation:
@@ 2.2 Iterative Markovian Fitting (IMF)
The SB solution is the unique path measure that is both Markovian and belongs to the reciprocal class of $\mathbb{Q}$ while matching marginals (leonard2014survey). This motivates the Iterative Markovian Fitting (IMF) algorithm (shi2023diffusionschrodingerbridgematching; Peluchetti2023DiffusionBM), which alternates between reciprocal and Markov projections:
$\mathbb{P}^{2n+1}=\operatorname{proj}_{\mathcal{M}}(\mathbb{P}^{2n}),\qquad\mathbb{P}^{2n+2}=\operatorname{proj}_{\mathcal{R}(\mathbb{Q})}(\mathbb{P}^{2n+1})$ (3)
These projections admit KL variational characterisations (A.1) and the iterations converge to $\mathbb{P}^{\star}$.
In practice, IMF is implemented by learning the drift of the Markovian projection via a bridge-matching loss (see A.1). Compared to Iterative Proportional Fitting (IPF), IMF preserves both marginals simultaneously and is more efficient (details in A.1).
@@ 3 Multi-Marginal temporal Schrödinger Bridge Matching
All proofs can be found in the Appendix A.6.
@@ 3.1 Multi-Marginal temporal Schrödinger Bridge Problem
Let $0=t_{0}<t_{1}<\cdots<t_{K}=T$ be a fixed time grid, and let $\mu_{0},...,\mu_{k},...,\mu_{T}\in\mathcal{P}(\mathbb{R}^{d})$ denote prescribed marginals at times $(t_{k})_{k=0,\dots,K}$, assuming $\mu_{t_{k}}\ll Q_{t_{k}}$ for all $k$. Given a reference process $\mathbb{Q}$ on $C([0,T],\mathbb{R}^{d})$, the multi-marginal Schrödinger Bridge problem (MMSB) is defined as
$\mathbb{P}^{\star}\;=\;\underset{\mathbb{P}\in\mathcal{P}(C)}{\operatorname{argmin}}\,\mathrm{KL}\!\left(\mathbb{P}\,\|\,\mathbb{Q}\right)\quad\text{subject to}\quad X_{t_{k}}\sim\mu_{k},\;\;k=0,\dots,K$ (4)
Note: Connection to multi-marginal Optimal Transport
If $\mathbb{Q}$ is associated with a Brownian motion, the induced reference coupling $\mathbb{Q}_{t_{0},\dots,t_{N}}$ is characterized by independent Gaussian increments $X_{t_{i+1}}-X_{t_{i}}\sim\mathcal{N}(0,\sigma^{2}(t_{i+1}-t_{i}))$. By evaluating the KL term, 4 can be rewritten as:
$\Pi^{\star}=\arg\min_{\Pi\in\mathcal{P}((\mathbb{R}^{d})^{N})}\Bigg\{\mathbb{E}_{X\sim\Pi}\Bigg[\sum_{i=0}^{K-1}\frac{1}{t_{i+1}-t_{i}}\,\|X_{t_{i+1}}-X_{t_{i}}\|^{2}\Bigg]-2\sigma^{2}T\mathcal{H}(\Pi)\;\;:\;\;\Pi_{i}=\mu_{t_{i}},\;\forall i\Bigg\}$
This is precisely an entropy-regularised multi-marginal OT problem with a time-structured quadratic cost $c(x_{0},\dots,x_{N})=\sum_{i=0}^{K-1}\frac{1}{t_{i+1}-t_{i}}\|x_{i+1}-x_{i}\|^{2}$ and entropy-regularisation parameter $\varepsilon=2\sigma^{2}$.
This formulation is particularly interesting when no better prior is available, and because of the clear interpretation it allows: when using a Brownian motion as prior, we are approaching quadratic OT. Note however that we do not rely on this assumption at all for theoretical results.
@@
These properties ensure that we can restrict our search to Markovian (and therefore reciprocal A.1) measures, which will be central to the projection algorithms introduced later. Finally, the explicit form of the solution further clarifies its structure.
@@
Under mild assumptions:
@@
where the $f_{i}$’s are functions of the Lagrange multipliers for the marginal constraints (see A.6.6).
@@ 3.2.1 Multi-marginal Markov and Reciprocal projections
By the Markovian projection theorem of gyongy1986, and as further developed in Peluchetti2023DiffusionBM; debortoli2021diffusionschrodingerbridgeapplications, the process $M^{\star}$ is Markov and matches the one-dimensional marginals of the original factorized law $\Pi$.
@@
Assume that $\sigma_{t}>0$. Let $M^{\star}=\operatorname{proj}_{\mathcal{M}}(\Pi)$ be the Markovian projection of $\Pi$ as in Definition $3.$ Then:
@@
$KL(\Pi\,\|\,M^{\star})=\frac{1}{2}\int_{0}^{T}\mathbb{E}_{\Pi_{t_{i},t}}\!\left[\frac{1}{\sigma_{t}^{2}}\Big\|\sigma_{t}^{2}\,\mathbb{E}_{\Pi_{t_{i+1}\mid t}}\!\big[\nabla\log\mathbb{Q}^{\,|t_{i},t_{i+1}}_{t}(X_{t_{i+1}}\mid X_{t})\,\big|\,X_{t},X_{t_{i}}\big]-v_{t}^{\star}(X_{t})\Big\|^{2}\right]dt$
@@
Conjecture 3.1 (Analogue of leonard2014survey Theorem 2.12).
@@ 3.2.2 Iterative Markovian Factorized Fitting
$KL(P^{n+1}\,\|\,P^{\star})\;\leq\;KL(P^{n}\,\|\,P^{\star})<\infty\quad\text{and}\quad\lim_{n\to\infty}KL(P^{n}\,\|\,P^{\star})=0$
@@
Under mild assumptions, the IMFF sequence $(\mathbb{P}^{n})_{n\in\mathbb{N}}$ admits at least one fixed point $\mathbb{P}^{\star}$, and we have:
@@
Moreover, denoting by $\mathbb{P}^{\mathrm{MMSB}}$ the solution of (MMSB) and by $\mathbb{P}^{\mathrm{pair}}$ the collage of pairwise Schrödinger Bridges, the limit of the IMFF sequence satisfies the inequality:
$KL(\mathbb{P}^{\mathrm{MMSB}}\,\|\,\mathbb{Q})\;\;=\;\;KL(\mathbb{P}^{\star}\,\|\,\mathbb{Q})\;\;\leq\;\;KL(\mathbb{P}^{\mathrm{pair}}\,\|\,\mathbb{Q})$
where $\mathbb{Q}$ is the chosen reference process. Thus, $\mathbb{P}^{\star}$ is the multi-marginal Schrödinger Bridge.